# Non-commutative rings with infinitely many maximal subrings

### v1

@@ 3. Rings with Infinitely Many Maximal Subrings
    for each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, the ring $\mathbb{I}(M)/M$ is a finite field. In particular, $C(T)/(C(T)\cap M)$ is a finite field.
@@
    for each $M\in Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)$, the center of $T/M$ is algebraically closed in $T/M$.
@@
    for each $M\in Max^{\prime}(T)$, there exist a natural number $n>1$ and a finite field $F$ such that $T/M\cong\mathbb{M}_{n}(F)$.
@@
    $Max^{\prime}(T)$ is finite and in fact $|Max^{\prime}(T)|\leq|RgMax(T)|$.
@@
    $Prm_{l}(T)\setminus Max_{l}(T)$, $Prm_{l}(T)\setminus Max(T)$, $Prm_{r}(T)\setminus Max_{r}(T)$ and $Prm_{r}(T)\setminus Max(T)$ are finite sets.
@@
    If $P$ a left (resp. right) primitive ideal of $T$, then either $P$ is a maximal ideal of $T$ or $P=\bigcap_{i\in I}M_{i}$, where $M_{i}\in Max_{l}(T)\cap Max(T)$, moreover in this case $I$ is infinite and $T/P$ is a domain (i.e., $P$ is a completely prime ideal of $T$). In particular, $T/P$ is embeddable in a division ring.
@@
$(1)$ for each maximal left ideal $A$ of $T$ which is not an ideal of $T$, by Theorem 2.7, $\mathbb{I}(A)$ is a maximal subring of $T$ which contains $A$. It is clear that whenever $A\neq B$ are in $Max_{l}^{\prime}(T)$, then $A+B=T$. Hence we conclude that $\mathbb{I}(A)\neq\mathbb{I}(B)$. Therefore the function $A\longmapsto\mathbb{I}(A)$, from $Max_{l}^{\prime}(T)$ in to $RgMax(T)$ is one-one and therefore $(1)$ holds. For $Max_{r}^{\prime}(T)$ the proof is similar.
@@
$(2)$ is an immediate consequence of Theorem 2.1 and $(1)$ (note that if $M\in Max_{l}^{\prime}(T)$, then $\mathbb{I}(M)/M\cong(End({}_{T}(T/M)))^{op}$ is a division ring). For the final part of $(2)$ note that $C(T)+M$ is a subring of $T$ which is contained in $\mathbb{I}(M)$ and therefore $(C(T)+M)/M$ is a subring of $\mathbb{I}(M)/M$. Now by the first part of $(2)$, we immediately deduce that $(C(T)+M)/M$ is a finite field and thus $C(T)/(C(T)\cap M)$ is a finite field too.
@@
$(3)$ if $T/M$ is a field (in particular, if $T/M$ is finite), then the conclusion is obvious. Hence assume that $T/M$ is an infinite division ring, then by Theorem 3.1, we deduce that the center of $T/M$ is algebraically closed in $T/M$ (for otherwise, $T/M$ and therefore $T$ have infinitely many maximal subrings, which is absurd).
@@
$(4)$ this is an immediate conclusion of Theorem 3.5 and the fact that $T/M$ has only finitely many maximal subrings (for $T$ has only finitely many maximal subrings).
@@
$(5)$ assume that $Max^{\prime}(T)=\{M_{i}\ |\ i\in I\}$, where $M_{i}\neq M_{j}$ for $i\neq j\in I$. Then by $(4)$, for each $i\in I$ there exist a natural number $n_{i}>1$ and a finite field $F_{i}$ such that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$. Therefore by $(4)$ of Theorem 2.7, we infer that for each $i\in I$, $M_{i}$ is contained in a maximal subring $S_{i}$ of $T$. Since for $i\neq j$, we have $M_{i}+M_{j}=T$, we deduce that $S_{i}\neq S_{j}$. Therefore $|I|\leq|RgMax(T)|$ and hence we are done.
@@
$(6)$ let $\{P_{i}\ |\ i\in I\}$ be a family of distinct left primitive ideals of $T$ such that for each $i\in I$, $P_{i}$ is not a maximal (left) ideal of $T$. Since for each $i\in I$, $P_{i}$ is a left primitive ideal of $T$, we deduce that there exists a maximal left ideal $M_{i}$ of $T$ such that $P_{i}=l.ann_{T}(T/M_{i})$. It is clear that whenever $i\neq j$, then $M_{i}\neq M_{j}$, for $P_{i}\neq P_{j}$. Also note that if $M_{i}$ is an ideal of $T$, then $P_{i}=M_{i}$ is a maximal (left) ideal of $T$ which is absurd. Thus for each $i\in I$ we deduce that $M_{i}\in Max_{l}^{\prime}(T)$ and therefore by $(1)$ we conclude that $I$ is finite. Thus the first and the second parts of $(6)$ hold. The proof for the next two parts are similar.
@@
$(7)$ let $P$ be a left primitive ideal of $T$. Hence $P$ is an intersection of a family of maximal left ideals of $T$. Since by $(1)$, $Max_{l}^{\prime}(T)$ is finite, we conclude that there exist $N_{1},\ldots,N_{k}\in Max_{l}^{\prime}(T)$, $k\geq 0$, and a family $\{M_{i}\ |\ i\in I\}\subseteq Max_{l}(T)\cap Max(T)$ such that $P=N_{1}\cap\cdots N_{k}\cap\bigcap_{i\in I}M_{i}$. By the comments preceding the proposition we infer that either $P=N_{1}\cap\cdots N_{k}$ or $P=\bigcap_{i\in I}M_{i}$. If $P=N_{1}\cap\cdots N_{k}$, then the ring $T/P$ embeds in the semisimple left $T$-module $T/N_{1}\times\cdots\times T/N_{k}$, as left $T$-module and therefore $T/P$ is a prime semisimple ring. Thus $T/P$ is a simple ring, i.e., $P$ is a maximal ideal of $T$. Hence assume that $P$ is not a maximal ideal of $T$ and $P=\bigcap_{i\in I}M_{i}$. Since $P$ is not maximal ideal of $T$, but is a prime ideal of $T$, we immediately conclude that $I$ is infinite. Finally note that for each $i\in I$, the ring $T/M_{i}$ is a division ring. Therefore $\prod_{i\in I}T/M_{i}$ is a reduced ring. Since $T/P$ embeds in $\prod_{i\in I}T/M_{i}$, we deduce that $T/P$ is a reduced ring. Thus $T/P$ is a reduced prime ring and hence is a domain. Therefore by [12, Theorem 1.2.4], $T/P$ is embeddable in a division ring.
@@
$(8)$ This is trivial by $(7)$ and the fact that $J(T)$ is the intersection of all left (resp. right) primitive ideals of $T$. ∎
@@
If $T$ is a simple ring then we are done by Theorem 3.5. Hence assume that $T$ is not a simple ring. If $T$ has only finitely many maximal subrings, then by $(7)$ of Proposition 3.10, we conclude that $T$ is embeddable in a division ring which is absurd. Thus $T$ has infinitely many maximal subrings. ∎
@@
Assume that $T$ has only finitely many maximal subrings, then by $(8)$ and $(5)$ of Proposition 3.10, we conclude that $0=J(T)=Q_{1}\cap\cdots\cap Q_{n}\cap N$, where $Q_{i}$ are maximal ideals of $T$ which are not maximal left ideals, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $Max(T)\cap Max_{l}(T)=\{M_{i}\ |\ i\in I\}$. Since $T$ is a prime ring, we infer that either there exists $k$ such that $Q_{k}=0$, i.e., $T$ is an infinite simple ring, or $N=0$. If $T$ is an infinite simple ring, then by hypothesise and Theorem 3.5, we deduce that $T$ has infinitely many maximal subrings which is absurd. Thus assume that $N=0$. Therefore $T$ embeds in $\prod_{i\in I}T/M_{i}$, and similar to the proof of $(7)$, we conclude that $T$ is embeddable in a division ring which is absurd. Thus $T$ has infinitely many maximal subrings. ∎
@@
Let $T$ be an infinite ring and $\{M_{i}\ |\ i\in I\}\subseteq Max_{l}(T)\cap Max(T)$ such that $\bigcap_{i\in I}M_{i}=0$. Then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$. In particular, if $T$ is a (left) quasi duo $J$-semisimple ring, then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$.
@@
Assume that $T$ has only finitely many maximal subrings. Hence by Theorem 3.1, for each $i\in I$, the center of $T/M_{i}$ is integrally closed in $T/M_{i}$, for otherwise if there exists $i\in I$ such that $T/M_{i}$ has a non-central element which is integral over the center of $T/M_{i}$, then $T/M_{i}$ and therefore $T$ have infinitely many maximal subrings, which is absurd. Now assume that $\alpha$ be a non-central element of $T$ which is integral over the center of $T$. Thus there exists $\beta\in T$ such that $\alpha\beta-\beta\alpha\neq 0$. Since $\bigcap_{i\in I}M_{i}=0$, we infer that there exists $i\in I$ such that $\alpha\beta-\beta\alpha\notin M_{i}$. Therefore $\alpha+M$ is a non-central element of the division ring $T/M_{i}$. Clearly, $\alpha+M_{i}$ is integral over the center of $T/M_{i}$, which is not possible. Thus we conclude that the center of $T$ is integrally closed in $T$. The final part is evident. ∎
Remark 3.14.
Let $T$ be a ring with only finitely many maximal subrings. Then $Max_{l}(T)$ is finite if and only if $T/J(T)$ is finite if and only if $Max_{r}(T)$ is finite. To see this, it suffices to show that if $Max_{l}(T)$ is finite, then $T/J(T)$ is finite. Since $Max_{l}(T)$ is finite, we immediately conclude that $T/J(T)$ is a semisimple ring. By assumption, $T$ and therefore $T/J(T)$ has only finitely many maximal subrings, therefore by Proposition 3.6, we deduce that each simple part of $T/J(T)$ is finite and therefore $T/J(T)$ is a finite ring.
For the next result, by the previous remark, if $T$ is a ring with only finitely many maximal subrings, then we may assume that $Max_{l}(T)$ is infinite and therefore by Proposition 3.10, we conclude that $Max_{l}(T)\cap Max(T)$ is infinite.
Corollary 3.15.
Let $T$ be an infinite ring with only finitely many maximal subrings. Then there exist finitely many distinct maximal ideals $Q_{1},\ldots,Q_{m}$, $m\geq 0$, such that:
  1. (1)
    for each $i$, there exist a natural number $n_{i}>1$ and a finite field $F_{i}$ such that $T/Q_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$.
  2. (2)
    $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, where $N=\bigcap_{M\in Max(T)\cap Max_{l}(T)}M$, where $N,Q_{1},\ldots,Q_{m}$ are incomparable.
  3. (3)
    $T/J(T)\cong R\times S$, where $R$ is a finite semisimple ring (in fact, $R=\prod_{i=1}^{n}\mathbb{M}_{n_{i}}(F_{i})$) and the center of $S$ is integrally closed in $S$.
Proof.
First note that by $(8)$ of Proposition 3.10, $J(T)=\bigcap_{M\in Max(T)}M$. Since by $(5)$ of Proposition 3.10, $Max^{\prime}(T)$ is finite we deduce that there exists finitely many distinct maximal ideals $Q_{1},\ldots,Q_{m}\in Max^{\prime}(T)$, $m\geq 0$, such that $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, where $N=\bigcap_{M\in Max(T)\cap Max_{l}(T)}M$ and clearly we may assume that $N$ and $Q_{i}$ are incomparable. Thus $(2)$ holds and note that $(1)$ is obvious by $(4)$ of Proposition 3.10. Since $Q_{i}$ are distinct maximal ideals which are incomparable to $N$, and $J(T)=Q_{1}\cap\cdots\cap Q_{m}\cap N$, we deduce the first part of $(3)$ (where $S=T/N$). The final part of $(3)$ is a consequence of Lemma 3.13. Hence we are done. ∎
@@
Let $T$ be a $K$-algebra over an infinite field $K$. Then either $T$ has infinitely many maximal subrings (in fact, $|RgMax(T)|\geq|K|$) or the following hold:
@@
    $T$ is a quasi duo ring. In particular, $Max(T)=Max_{l}(T)=Max_{r}(T)=Prm_{l}(T)=Prm_{r}(T)$.
@@
    For each $M\in Max(T)$, the center of $T/M$ is integrally closed in $T/M$.
@@
    $J(T)=\bigcap_{M\in Max(T)}M$ and $T/J(T)$ is a reduced ring. In particular, $N(T)\subseteq J(T)$. Moreover, the center of $T/J(T)$ is integrally closed in $T/J(T)$.
@@
If $T$ has a maximal left ideal $M$, which is not an ideal of $T$, then by Corollary 2.2, we deduce that $|K|\leq|Max_{l}(T)\setminus Max(T)|$. Thus by $(1)$ of Proposition 3.10, we conclude that $|K|\leq|RgMax(T)|$. Hence if $|RgMax(T)|<|K|$ we deduce that each maximal left ideal of $T$ is an ideal of $T$, i.e., $T$ is a left quasi duo ring, and by a similar proof $T$ is a right quasi duo ring. Hence $T$ is a quasi duo ring. Thus $(1)$ holds. $(2)$ is an immediate consequence of Proposition 3.1. For $(3)$, by $(1)$ we infer that $J(T)$ is the intersection of all maximal ideals of $T$ and therefore $T/J(T)$ embeds in the reduce ring $\prod_{M\in Max(T)}T/M$ (note that $T$ is a quasi duo ring and therefore $T/M$ is a division ring for each $M\in Max(T)$). Hence $T/J(T)$ is a reduced ring and thus $N(T)\subseteq J(T)$. The final part of $(3)$ is obvious by Lemma 3.13 (note, $T/J(T)$ is an infinite ring, for it contains a copy of the infinite field $K$). ∎
@@
Let $T$ be a prime $K$-algebra over an infinite field $K$ with $J(T)=0$, then either $|RgMax(T)|\geq|K|$ or $T$ embeds in a division ring (therefore $T$ is a domain).
@@
Assume that $|RgMax(T)|<|K|$, then by the proof of the previous theorem $T$ embeds in the reduced ring $R:=\prod_{M\in Max(T)}T/M$. Thus $T$ is a reduce prime ring. Therefore $T$ is a domain. Since for each maximal ideal $M$ of $T$, $T/M$ is a division ring, we infer that $T$ is a strongly von Neumann regular ring. Hence by [12, Theorem 1.2.4], we conclude that $T$ is embeddable in a division ring (note, in fact $T$ is a division ring, for $T$ is a strongly von Neumann regular domain and therefore $C(T)$ is integrally closed in $T$, by Proposition 3.1). ∎
If $R$ is a commutative ring and $M$ be a maximal ideal of $R$ such that $Char(R/M)=0$, then $R/M$ is a field with infinitely many maximal subrings, see [10, Corollary 1.5]. Therefore $R$ has infinitely many maximal subrings. For non-commutative rings we have the following result.
@@
    For each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
@@
    For each $M\in Max^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
@@
    If $P$ is a left/right primitive ideal of $T$, then either $P$ is a maximal left and right ideal of $T$ or $Char(T/P)$ is a prime number.

### v2

@@ 3. Rings with Infinitely Many Maximal Subrings
    For each $M\in\operatorname{{Max}_{\ell}}^{\prime}(T)\cup\operatorname{{Max}_{r}}^{\prime}(T)$, the ring $\mathbb{I}(M)/M$ is a finite field. In particular, $C(T)/(C(T)\cap M)$ is a finite field.
@@
    For each $M\in\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)=\operatorname{{Max}_{r}}(T)\cap\operatorname{Max}(T)$, the center of $T/M$ is algebraically closed in $T/M$.
@@
    For each $M\in\operatorname{Max}^{\prime}(T)$, there exist a natural number $n>1$ and a finite field $F$ such that $T/M\cong\mathbb{M}_{n}(F)$.
@@
    $\operatorname{Max}^{\prime}(T)$ is finite and in fact $|\operatorname{Max}^{\prime}(T)|\leq|\operatorname{RgMax}(T)|$.
@@
    $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{{Max}_{\ell}}(T)$, $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{Max}(T)$, $\operatorname{{Prm}_{r}}(T)\setminus\operatorname{{Max}_{r}}(T)$ and $\operatorname{{Prm}_{r}}(T)\setminus\operatorname{Max}(T)$ are finite sets.
@@
    If $P$ a left (resp. right) primitive ideal of $T$, then either $P$ is a maximal ideal of $T$ or $P=\bigcap_{i\in I}M_{i}$, where $M_{i}\in\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)$, moreover in this case $I$ is infinite and $T/P$ is a domain (i.e., $P$ is a completely prime ideal of $T$). In particular, $T/P$ is embeddable in a division ring.
@@
$(\bf 1)$ For each maximal left ideal $A$ of $T$ which is not an ideal of $T$, by Theorem 2.7, $\mathbb{I}(A)$ is a maximal subring of $T$ which contains $A$. It is clear that whenever $A\neq B$ are in $\operatorname{{Max}_{\ell}}^{\prime}(T)$, then $A+B=T$ and consequently $\mathbb{I}(A)\neq\mathbb{I}(B)$. Therefore the function $A\longmapsto\mathbb{I}(A)$, from $\operatorname{{Max}_{\ell}}^{\prime}(T)$ into $\operatorname{RgMax}(T)$ is one-one, which complete the proof of $(1)$ for $\operatorname{{Max}_{\ell}}(T)$. The proof for $\operatorname{{Max}_{r}}^{\prime}(T)$ is analogous.
@@
$(\bf 2)$ The first part is an immediate consequence of Theorem 2.1 and $(1)$ (note that if $M\in\operatorname{{Max}_{\ell}}^{\prime}(T)$, then $\mathbb{I}(M)/M\cong(\operatorname{End}({}_{T}(T/M)))^{op}$ is a division ring). For the last part of $(2)$, obviously $C(T)+M$ is a subring of $T$ which is contained in $\mathbb{I}(M)$. It follows that $(C(T)+M)/M$ is a subring of $\mathbb{I}(M)/M$. Now by the first part of $(2)$, we deduce that $(C(T)+M)/M$ is a finite field. A fortiori, $C(T)/(C(T)\cap M)$ is a finite field.
@@
$(\bf 3)$ If $T/M$ is a field (in particular, if $T/M$ is finite), then the conclusion is obvious. Hence assume that $T/M$ is an infinite division ring. By Theorem 3.1, we deduce that the center of $T/M$ is algebraically closed in $T/M$, because otherwise, $T/M$ and therefore $T$ have infinitely many maximal subrings, which is absurd.
@@
$(\bf 4)$ This is an immediate consequence of Theorem 3.5 and the fact that $T/M$ has only finitely many maximal subrings (note that $T$ has only finitely many maximal subrings and thus $T/M$ has only finitely many maximal subrings).
@@
$(\bf 5)$ Suppose that $\operatorname{Max}^{\prime}(T)=\{M_{i}\ |\ i\in I\}$, where $M_{i}\neq M_{j}$ for $i\neq j\in I$. Then by $(4)$, for each $i\in I$ there exist a natural number $n_{i}>1$ and a finite field $F_{i}$ such that $T/M_{i}\cong\mathbb{M}_{n_{i}}(F_{i})$. Therefore by Theorem 2.7(4)(4), we infer that for each $i\in I$, $M_{i}$ is contained in a maximal subring $S_{i}$ of $T$. Since for $i\neq j$, we have $M_{i}+M_{j}=T$, we deduce that $S_{i}\neq S_{j}$. Therefore $|I|\leq|\operatorname{RgMax}(T)|$ and hence we are done.
@@
$(\bf 6)$ Let $\{P_{i}\ |\ i\in I\}$ be a family of distinct left primitive ideals of $T$ such that for any $i\in I$, $P_{i}$ is not a maximal (left) ideal of $T$. Since for each $i\in I$, $P_{i}$ is a left primitive ideal of $T$, we deduce that there exists a maximal left ideal $M_{i}$ of $T$ such that $P_{i}=\operatorname{l.ann}_{T}(T/M_{i})$. It is clear that whenever $i\neq j$, then $M_{i}\neq M_{j}$, for $P_{i}\neq P_{j}$. Also note that if $M_{i}$ is an ideal of $T$, then $P_{i}=M_{i}$ is a maximal (left) ideal of $T$ which is absurd. Thus for each $i\in I$ we obtain that $M_{i}\in\operatorname{{Max}_{\ell}}^{\prime}(T)$ and therefore by $(1)$ we conclude that $I$ is finite. It follows that $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{{Max}_{\ell}}(T)$, $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{Max}(T)$ are finite. The proof for $\operatorname{{Prm}_{r}}(T)\setminus\operatorname{{Max}_{r}}(T)$ and $\operatorname{{Prm}_{r}}(T)\setminus\operatorname{Max}(T)$ are analogous.
@@
$(\bf 7)$ Let $P$ be a left primitive ideal of $T$. Hence $P$ is an intersection of a family of maximal left ideals of $T$. Since by $(1)$, $\operatorname{{Max}_{\ell}}^{\prime}(T)$ is finite, we conclude that there exist $N_{1},\ldots,N_{k}\in\operatorname{{Max}_{\ell}}^{\prime}(T)$, $k\geq 0$, and a family $\{M_{i}\ |\ i\in I\}\subseteq\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)$ such that $P=N_{1}\cap\cdots N_{k}\cap N$, where $N=\bigcap_{i\in I}M_{i}$. By the comments preceding the proposition we infer that either $P=N_{1}\cap\cdots N_{k}$ or $P=\bigcap_{i\in I}M_{i}$. If $P=N_{1}\cap\cdots N_{k}$, then the ring $T/P$ embeds in the semisimple left $T$-module $T/N_{1}\times\cdots\times T/N_{k}$, as left $T$-modules. This implies that $T/P$ is a prime semisimple ring. Consequently $T/P$ is a simple ring, i.e., $P$ is a maximal ideal of $T$. Hence we may suppose that $P$ is not a maximal ideal of $T$ and $P=N=\bigcap_{i\in I}M_{i}$. Since $P$ is not maximal ideal of $T$, but is a prime ideal of $T$, we conclude that $I$ is infinite. Finally note that for each $i\in I$, the ring $T/M_{i}$ is a division ring. A fortiori, $\prod_{i\in I}T/M_{i}$ is a reduced ring. Since $T/P$ embeds in $\prod_{i\in I}T/M_{i}$, we deduce that $T/P$ is a reduced ring. Thus $T/P$ is a reduced prime ring and consequently $T/P$ is a domain. It follows by [12, Theorem 1.2.4] that $T/P$ is embeddable in a division ring.
@@
$(\bf 8)$ This is trivial by $(7)$ and the fact that $J(T)$ is the intersection of all left (resp. right) primitive ideals of $T$. ∎
The previous observations yield the following quick corollaries.
@@
If $T$ is a simple ring then we are done by Theorem 3.5. Hence assume that $T$ is not a simple ring. If $T$ has only finitely many maximal subrings, then by Proposition 3.10(7)(7), we conclude that $T$ is embeddable in a division ring which is absurd. Thus $T$ has infinitely many maximal subrings. ∎
@@
Assume that $T$ has only finitely many maximal subrings, then by $(8)$ and $(5)$ of Proposition 3.10, we conclude that $0=J(T)=Q_{1}\cap\cdots\cap Q_{n}\cap N$, where $Q_{i}$ are maximal ideals of $T$ which are not maximal left ideals, $n\geq 0$ and $N=\bigcap_{i\in I}M_{i}$, where $\operatorname{Max}(T)\cap\operatorname{{Max}_{\ell}}(T)=\{M_{i}\ |\ i\in I\}$. Since $T$ is a prime ring, we infer that either there exists $k$ such that $Q_{k}=0$, i.e., $T$ is an infinite simple ring, or $N=0$. If $T$ is an infinite simple ring, then by hypothesise and Theorem 3.5, we deduce that $T$ has infinitely many maximal subrings which is absurd. Thus assume that $N=0$. Therefore $T$ embeds in $\prod_{i\in I}T/M_{i}$, and similar to the proof of Proposition 3.10(7)(7), we conclude that $T$ is embeddable in a division ring which is impossible. Thus $T$ has infinitely many maximal subrings. ∎
@@
Let $T$ be an infinite ring and $\bigcap_{i\in I}M_{i}=0$ where $\{M_{i}\ |\ i\in I\}\subseteq\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)$. Then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$. In particular, if $T$ is a (left) quasi-duo $J$-semisimple ring, then either $T$ has infinitely many maximal subring or $C(T)$ is integrally closed in $T$.
@@
Assume that $T$ has only finitely many maximal subrings. Consequently for any $i\in I$, the division ring $T/M_{i}$ has only finitely many maximal subrings. Hence by Theorem 3.1, for each $i\in I$, the center of $T/M_{i}$ is integrally closed in $T/M_{i}$, for otherwise if there exists $i\in I$ such that $T/M_{i}$ has a non-central element which is integral over the center of $T/M_{i}$, then $T/M_{i}$ has infinitely many maximal subrings, which is absurd. Now assume that $\alpha$ is a non-central element of $T$ which is integral over the center of $T$. Thus there exists $\beta\in T$ such that $\alpha\beta-\beta\alpha\neq 0$. Since $\bigcap_{i\in I}M_{i}=0$, we infer that there exists $i\in I$ such that $\alpha\beta-\beta\alpha\notin M_{i}$. Therefore $\alpha+M$ is a non-central element of the division ring $T/M_{i}$. Clearly, $\alpha+M_{i}$ is integral over the center of $T/M_{i}$, which is impossible. Thus, the center of $T$ is integrally closed in $T$. The last part is evident. ∎
@@
Let $T$ be a $K$-algebra over an infinite field $K$. Then either $T$ has infinitely many maximal subrings (in fact, $|\operatorname{RgMax}(T)|\geq|K|$) or the following hold:
@@
    $T$ is a quasi-duo ring. In particular, $\operatorname{Max}(T)=\operatorname{{Max}_{\ell}}(T)=\operatorname{{Max}_{r}}(T)=\operatorname{{Prm}_{\ell}}(T)=\operatorname{{Prm}_{r}}(T)$.
@@
    For each $M\in\operatorname{Max}(T)$, the center of $T/M$ is integrally closed in $T/M$.
@@
    $J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$ and $T/J(T)$ is a reduced ring. In particular, $N(T)\subseteq J(T)$. Moreover, the center of $T/J(T)$ is integrally closed in $T/J(T)$.
@@
If $T$ has a maximal left ideal $M$, which is not an ideal of $T$, then by Corollary 2.2, we deduce that $|K|\leq|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|$. Thus by Proposition 3.10(1)(1), we conclude that $|K|\leq|\operatorname{RgMax}(T)|$. Hence if $|\operatorname{RgMax}(T)|<|K|$ we obtain that each maximal left ideal of $T$ is an ideal of $T$, i.e., $T$ is a left quasi-duo ring, and by an analogous proof, $T$ is a right quasi-duo ring. Consequently $T$ is a quasi-duo ring. Thus $(1)$ holds. $(2)$ is an immediate consequence of Proposition 3.1. For $(3)$, by $(1)$ we infer that $J(T)$ is the intersection of all maximal ideals of $T$ and therefore $T/J(T)$ embeds in the reduce ring $\prod_{M\in\operatorname{Max}(T)}T/M$ (note that $T$ is a quasi-duo ring and therefore $T/M$ is a division ring for each $M\in\operatorname{Max}(T)$). Hence $T/J(T)$ is a reduced ring and thus $N(T)\subseteq J(T)$. The last part of $(3)$ is obvious by Lemma 3.13 (note, $T/J(T)$ is an infinite ring, because it contains a copy of the infinite field $K$). ∎
@@
Let $T$ be a prime $K$-algebra over an infinite field $K$ with $J(T)=0$, then either $|\operatorname{RgMax}(T)|\geq|K|$ or $T$ embeds in a division ring; in particular, $T$ is a domain.
@@
Assume that $|\operatorname{RgMax}(T)|<|K|$, then by the proof of the previous theorem $T$ embeds in the reduced ring $R:=\prod_{M\in\operatorname{Max}(T)}T/M$. Thus $T$ is a reduce prime ring. Therefore $T$ is a domain. Since for each maximal ideal $M$ of $T$, $T/M$ is a division ring, we infer that $T$ is a strongly von Neumann regular ring. Hence by [12, Theorem 1.2.4], we conclude that $T$ is embeddable in a division ring, as desired.∎
If $R$ is a commutative ring and $M$ is a maximal ideal of $R$ such that $\operatorname{Char}(R/M)=0$, then $R/M$ is a field with infinitely many maximal subrings, by [10, Corollary 1.5]. Consequently, $R$ itself has infinitely many maximal subrings. For noncommutative rings we obtain the following analogue.
@@
    For each $M\in\operatorname{{Max}_{\ell}}^{\prime}(T)\cup\operatorname{{Max}_{r}}^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
@@
    For each $M\in\operatorname{Max}^{\prime}(T)$, there exists a prime number $p$ such that $pT\subseteq M$.
@@
    If $P$ is a left/right primitive ideal of $T$, then either $T/P$ is a division ring or $\operatorname{Char}(T/P)$ is a prime number.